OpenStudy (anonymous):

Medal WILL be rewarded! SHOW WORK!!! Find the value of x to the nearest degree.

OpenStudy (usarmy3947):

ill help

OpenStudy (anonymous):

83 23 53 67

OpenStudy (anonymous):

wow okay :)

OpenStudy (anonymous):

Which trig ratio deals with the opposite side and the hypotenuse? Hint: Think about SOH-CAH-TOA.. @jbell0

OpenStudy (turingtest):

what are the definitions of the trigonometric functions?

OpenStudy (anonymous):

YES THE GREAT CHIEF SOHCAHTOA!!! sin, opposite/hypotenuse cos, adjacent/hypotenuse tan, opposite/adjacent

OpenStudy (turingtest):

so which of the side lengths are given in the drawing?

OpenStudy (usarmy3947):

wait here

OpenStudy (anonymous):

the opposite and the hypotenuse...so it would be sin?

OpenStudy (usarmy3947):

sin(?)

OpenStudy (turingtest):

exactly :) so how can we get the angle from that?

OpenStudy (usarmy3947):

ok whats the ?

OpenStudy (usarmy3947):

its a letter in your question u left alone

OpenStudy (usarmy3947):

sorry some is reporting me for giving u the answer

OpenStudy (anonymous):

sooooo......x=sin-1(opp/hyp)

OpenStudy (usarmy3947):

here look

OpenStudy (turingtest):

yep :)

OpenStudy (usarmy3947):

sin(x) = 3/ √58 =0.393919 radian x= arcsin(0.393919) = ?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

thanks you guys

OpenStudy (turingtest):

very welcome

OpenStudy (usarmy3947):

ur welcome and that the right answer